Suppose the United States does not produce any baseball hats domestically but imports them from foreign producers. Initially, demand is Q 2000-5p and supply (from foreign producers) is Qs- 400+3p. Determine the equilibrium price and quantity. The government then decides that no more than 500 baseball hats should be imported per period and imposes a quota at that level. How does this quota affect the equilibrium price and quantity? Show the solution using a graph and calculate the numerical answer. How might this quota affect the market for cowboy hats (a substitute good)? Please show all work and explain your answers 2.
Answer #1

Solution :- At an equilibrium, Demand = Supply.

2000 - 5P = 400 + 3P

2000 - 400 = 3P + 5P

1600 = 8P

P = 1600 / 8

P = $ 200

Put the value of P = 200 in the demand equation.

Q = 2000 - 5 * 200

Q = 2000 - 1000

Q = 1000.

Conclusion :-

Equilibrium price (P) $ 200
Equilibrium quantity (Q) 1000.